A company manufactures fuses. The percentage of non-defective fuses is 95.4%. A sample of 9 fuse was selected. Calculate the probability of selecting at least 3 defective fuses.   


1
Expert's answer
2021-07-13T13:04:56-0400

The Poisson distribution is

"P\left(X=x\right)=\frac{e^{-\lambda }\lambda ^x}{x!}"

Let "x" denote the defective fuse

"P=100\%-95.4\%=4.6\%=0.046"

"n=9"

"mean=\lambda =np=9\times 0.046=0.414"

"P\left(at\,least\:3\:defective\:fuses\right)=\sum_{x \ge 3} \frac{e^{-0.414}0.414^3}{3}=0.0156"